# An Analysis of LIGO Glitches Using t-SNE During the First Part of the Fourth LIGO-Virgo-KAGRA Observing Run

### v1

@@ 3 Glitches during O4a at LHO
The rate of glitches with an SNR greater than 10 was significantly lower at LHO during O4a (approximately 6 per hour) compared to O3 (around 17 per hour). However, LHO was notably affected by glitches with lower SNR. When considering glitches with an SNR exceeding 6.5, Hanford experienced an average of 77 glitches per hour during O4a — more than double the glitch rate observed during O3, which was approximately 33 per hour [soni2025ligo]. This increase is attributed to the presence of broadband transient noise between 20–40 Hz during O4a, which occasionally extended to peak frequencies around 50 Hz. The predominance of this frequency range is evident in Figure 13, which shows the histogram of glitches during O4a with an SNR greater than 6.
@@
In June, the bottom-center region of the t-SNE map becomes increasingly prominent, persists through July, and decreases in August and September, coinciding with the activation of other groups and resulting in a more uniform overall distribution. This decrease was driven by the implementation of an improved alignment control system [71927], which substantially reduced the occurrence of glitches with SNR greater than 7.5 [alog71933]. This is also reflected in Figure 15, which shows a reduction in the glitch rate from 138.7 to 68.1 glitches per hour between July and August.
From October through January, the large cluster H1, together with H2, became the most persistent feature in the t-SNE projection, collectively accounting for at least 70% of the observed glitches. Although the implementation of the alignment controller substantially contributed to glitch mitigation, an increase in broadband glitches emerged in October and persisted throughout the remainder of the observing run, leading to the predominance of these two clusters. This increase coincided with the relocation of the LSC (Length Sensing and Control) feedforward path from the ETMX (End Test Mass of the X-arm) to the ETMY actuator [H1alog73420], which initially reduced noise above $50\text{\,}\mathrm{H}\mathrm{z}$ but introduced additional noise at lower frequencies. Subsequent investigations suggested that the ETMY R0  [^1] R0 refers to the uppermost stage of the reaction chain [colgan2023detecting]. actuator, responsible for length control, may not have been operating properly at that time [H1alog73447], potentially generating excess scattered light and contributing to the observed rise in glitch rate.
A faulty cable connected to the ETMY R0 actuator was identified as the source of the malfunction [alog73522]; its replacement led to a slight decrease in the glitch rate. This was followed by a new increase, possibly associated with the implementation of an updated feedforward filter in the Signal Recycling Cavity Length (SRCL) control loop of the Length Sensing and Control (LSC) system [alog73561], which was later revised, resulting in a marked decrease in the glitch rate from October 23 onward [alog73662] — this decrease is highlighted by the vertical pink dashed line in Figure 18, where the dark red vertical bars represent the evolution of the hourly glitch rate throughout the entire O4a.
@@
A bicoherence  [^2] Bicoherence [choudhury2008bispectrum] is a diagnostic tool that assesses nonlinear coupling between two frequencies. If two frequencies, $f_{1}$ and $f_{2}$, are present in a signal, bicoherence can indicate whether a new frequency component at $f_{1}+f_{2}$ (or $f_{1}-f_{2}$) is being generated through nonlinear interactions. analysis revealed that these broadband glitches were related to the electrostatic drive (ESD) system in the detector [alog71005]. Positioned behind the test mass, the ESD applied controlled forces to the mirror to maintain the interferometer in its locked state — that is, when all the optical cavities are held on resonance through feedback control — enabling it to operate in observation mode for gravitational-wave detection. The ESD consists of a constant voltage component — commonly referred to as the bias — and multiple lower-voltage components applied at four sections: upper left (UL), upper right (UR), lower left (LL), and lower right (LR). Each of these contributes independently, with voltages that typically vary in amplitude according to the control needs. Together, they enable fine adjustments to the mirror’s position along the length degrees of freedom.
Along with the hourly glitch rate during O4a, Figure 18 also presents the RMS value of the ESD voltage for the UL component at ETMX (in black). The impact of the new alignment control system is evident at the beginning of August, with a significant drop in both ESD RMS and glitch rate. Another key moment occurs in October, when both show a substantial increase, related to the reallocation of the LSC feedforward from ETMX to ETMY. The main purpose of this plot is to illustrate that significant changes in the ESD indeed affected the rate of the broadband glitches at LHO during O4a. Although other factors also contributed to transient noise in the detector and influenced the curve behavior [alog73550, alog73740], a clear correlation can be observed between variations in the ESD RMS signal and fluctuations in the glitch rate, supporting the hypothesis of nonlinear coupling between the ESD actuation system and the detector strain output.
@@
The correlation between the glitch rate and ESD RMS for these two groups was also quantitatively estimated using the Pearson correlation coefficient (rr[benesty2009pearson]. The analysis confirmed a lack of correlation for Group $H_{4}$ ($r=0.12$) and a strong positive correlation for Group $H_{1}$ with $r=0.76$. For these plots and numbers, we focus on data collected after October, when the ESD exhibited significant oscillations. Between August and October, the correlation was not evident, as the ESD remained nearly constant (see Figure 18). Likewise, the same plot using data prior to August does not show this correlation, which remains not fully understood, given that broadband glitches appear to originate from the same key source.
@@
Although these glitches affected the entire O4a, dedicated testing and efforts by the LHO team led to the implementation of a new feedback control configuration, which significantly reduced the noise and resulted in a more stable and stationary system by lowering the ESD drive voltage. This improved configuration was adopted for the O4b observing run, which began with no recurrence of these glitches [alog74887, alog74939, alog76449, alog76459].
@@ 4 Conclusions and Discussions
The technique aims to uncover the monthly behavior of glitches in gravitational-wave detectors such as LIGO, Virgo, and KAGRA. The only requirement is access to Omicron data — or a similar tool capable of identifying triggers that encode glitch morphology. Once the glitches are tracked, the ideal next step — as demonstrated in this study — is to examine relevant detector logbook entries that document changes or interventions at the observatories, in order to correlate the behavior of glitch groups with those modifications. By analyzing monthly trends during O4a, we were able to correlate certain glitch groups with environmental or instrumental conditions, providing strong indications of their possible origin. As discussed in the methods paper [ferreira2024using], this approach can be applied at different timescales depending on the specific goals. Although the methodology autonomously identifies glitch groups, human interpretation — along with association to auxiliary channels and documented changes in the observatories — remains essential to understanding their physical origin.

### v2

@@ 3 Glitches during O4a at LHO
The rate of glitches with an SNR greater than 10 was significantly lower at LHO during O4a (approximately 6 per hour) compared to O3 (around 17 per hour). However, LHO was notably affected by glitches with lower SNR. When considering glitches with an SNR exceeding 6.5, Hanford experienced an average of 77 glitches per hour during O4a — more than double the glitch rate observed during O3, which was approximately 33 per hour [7]. This increase is attributed to the presence of broadband transient noise between 20–40 Hz during O4a, which occasionally extended to peak frequencies around 50 Hz. The predominance of this frequency range is evident in Figure 13, which shows the histogram of glitches during O4a with an SNR greater than 6.
@@
In June, the bottom-center region of the t-SNE map becomes increasingly prominent, persists through July, and decreases in August and September, coinciding with the activation of other groups and resulting in a more uniform overall distribution. This decrease was driven by the implementation of an improved alignment control system [55], which substantially reduced the occurrence of glitches with SNR greater than 7.5 [56]. This is also reflected in Figure 15, which shows a reduction in the glitch rate from 138.7 to 68.1 glitches per hour between July and August.
From October through January, the large cluster H1, together with H2, became the most persistent feature in the t-SNE projection, collectively accounting for at least 70% of the observed glitches. Although the implementation of the alignment controller substantially contributed to glitch mitigation, an increase in broadband glitches emerged in October and persisted throughout the remainder of the observing run, leading to the predominance of these two clusters. This increase coincided with the relocation of the LSC (Length Sensing and Control) feedforward path from the ETMX (End Test Mass of the X-arm) to the ETMY actuator [57], which initially reduced noise above $50\text{\,}\mathrm{H}\mathrm{z}$ but introduced additional noise at lower frequencies. Subsequent investigations suggested that the ETMY R0  [^1] R0 refers to the uppermost stage of the reaction chain [58]. actuator, responsible for length control, may not have been operating properly at that time [59], potentially generating excess scattered light and contributing to the observed rise in glitch rate.
A faulty cable connected to the ETMY R0 actuator was identified as the source of the malfunction [60]; its replacement led to a slight decrease in the glitch rate. This was followed by a new increase, possibly associated with the implementation of an updated feedforward filter in the Signal Recycling Cavity Length (SRCL) control loop of the Length Sensing and Control (LSC) system [61], which was later revised, resulting in a marked decrease in the glitch rate from October 23 onward [62] — this decrease is highlighted by the vertical pink dashed line in Figure 18, where the dark red vertical bars represent the evolution of the hourly glitch rate throughout the entire O4a.
@@
A bicoherence  [^2] Bicoherence [63] is a diagnostic tool that assesses nonlinear coupling between two frequencies. If two frequencies, $f_{1}$ and $f_{2}$, are present in a signal, bicoherence can indicate whether a new frequency component at $f_{1}+f_{2}$ (or $f_{1}-f_{2}$) is being generated through nonlinear interactions. analysis revealed that these broadband glitches were related to the electrostatic drive (ESD) system in the detector [64]. Positioned behind the test mass, the ESD applied controlled forces to the mirror to maintain the interferometer in its locked state — that is, when all the optical cavities are held on resonance through feedback control — enabling it to operate in observation mode for gravitational-wave detection. The ESD consists of a constant voltage component — commonly referred to as the bias — and multiple lower-voltage components applied at four sections: upper left (UL), upper right (UR), lower left (LL), and lower right (LR). Each of these contributes independently, with voltages that typically vary in amplitude according to the control needs. Together, they enable fine adjustments to the mirror’s position along the length degrees of freedom.
Along with the hourly glitch rate during O4a, Figure 18 also presents the RMS value of the ESD voltage for the UL component at ETMX (in black). The impact of the new alignment control system is evident at the beginning of August, with a significant drop in both ESD RMS and glitch rate. Another key moment occurs in October, when both show a substantial increase, related to the reallocation of the LSC feedforward from ETMX to ETMY. The main purpose of this plot is to illustrate that significant changes in the ESD indeed affected the rate of the broadband glitches at LHO during O4a. Although other factors also contributed to transient noise in the detector and influenced the curve behavior [65, 66], a clear correlation can be observed between variations in the ESD RMS signal and fluctuations in the glitch rate, supporting the hypothesis of nonlinear coupling between the ESD actuation system and the detector strain output.
@@
The correlation between the glitch rate and ESD RMS for these two groups was also quantitatively estimated using the Pearson correlation coefficient (rr[54]. The analysis confirmed a lack of correlation for Group $H_{4}$ ($r=0.12$) and a strong positive correlation for Group $H_{1}$ with $r=0.76$. For these plots and numbers, we focus on data collected after October, when the ESD exhibited significant oscillations. Between August and October, the correlation was not evident, as the ESD remained nearly constant (see Figure 18). Likewise, the same plot using data prior to August does not show this correlation, which remains not fully understood, given that broadband glitches appear to originate from the same key source.
@@
Although these glitches affected the entire O4a, dedicated testing and efforts by the LHO team led to the implementation of a new feedback control configuration, which significantly reduced the noise and resulted in a more stable and stationary system by lowering the ESD drive voltage. This improved configuration was adopted for the O4b observing run, which began with no recurrence of these glitches [67, 68, 69, 70].
@@ 4 Conclusions and Discussions
The technique aims to uncover the monthly behavior of glitches in gravitational-wave detectors such as LIGO, Virgo, and KAGRA. The only requirement is access to Omicron data — or a similar tool capable of identifying triggers that encode glitch morphology. Once the glitches are tracked, the ideal next step — as demonstrated in this study — is to examine relevant detector logbook entries that document changes or interventions at the observatories, in order to correlate the behavior of glitch groups with those modifications. By analyzing monthly trends during O4a, we were able to correlate certain glitch groups with environmental or instrumental conditions, providing strong indications of their possible origin. As discussed in the methods paper [35], this approach can be applied at different timescales depending on the specific goals. Although the methodology autonomously identifies glitch groups, human interpretation — along with association to auxiliary channels and documented changes in the observatories — remains essential to understanding their physical origin.